### v1

@@ 1 Introduction
Policy-gradient (PG) methods are a widely used class of reinforcement learning (RL) algorithms that optimize a parameterized policy directly from sampled trajectories. In LLM post-training, PG-style algorithms are now standard for improving reasoning behavior, with widely used objectives including REINFORCE Williams (1992), Proximal Policy Optimization (PPO) Schulman et al. (2017), and GRPO Shao et al. (2024).
These methods are typically implemented in an on-policy or near on-policy regime, where data generation and learning must occur sequentially after each other. This coupling becomes a major systems bottleneck for long-sequence reasoning: rollout lengths are highly variable and exhibit a long tail Hu et al. (2025)., creating pipeline bubbles and poor compute utilization in synchronous RL training.
To improve throughput, several recent systems decouple rollout generation from learning via asynchronous, pipelined training. Prior works such as asynchronous RLHF Noukhovitch et al. (2024) AReaL Fu et al. (2025), LlamaRL Wu et al. (2025), and PipelineRL Piché et al. (2025) demonstrate that overlapping generation with optimization can substantially reduce end-to-end wall-clock time. However, this efficiency comes at the cost of training quality: the sampler generates trajectories under a lagging behavior policy while the learner continues to update, making training effectively off-policy which leads to learning degradation or even collapse.
In this work, we focus on the optimization failure mode created by off-policy training and propose variance-targeted controls that can be applied on top of existing asynchronous RL frameworks. We find that the underlying mechanism behind asynchronous training instabilities is higher-variance gradient estimates. We show that the effective sample size Kish (1995) (ESS) – a well-known diagnostic for importance sampled estimates – collapses when the sampler and learner distributions significantly diverge and that this collapse predicts gradient volatility and KL spikes across reasoning tasks. Motivated by this diagnosis, we propose Variance Controlled Policy Optimization (VCPO), a simple drop-in method that directly controls the variance of the policy-gradient estimator with two complementary components: (i) an effective sample size guided step scaling that down-weights unreliable updates and (ii) a closed-form minimum-variance off-policy reward baseline based on per-trajectory gradient norms, avoiding training an auxiliary critic or value model. To make (ii) practical at scale, we present the first scalable implementation for exact per-example gradient-norm computation with negligible runtime overhead, and leverage these statistics to build a low-overhead, gradient-aware baseline for variance reduction.
@@ 2 Background
2.1 Reinforcement Learning for Large Language Models
@@
2.2 The Role of Effective Sample Size in Off-Policy RL
@@
2.3 RL stabilization methods
While prior methods have been proposed to stabilize LLM RL in on-policy or near-on-policy settings, we discuss these works as they closely related to asynchronous training’s highly off-policy issues. These approaches broadly fall into three categories: (1) masking/clipping to suppress outlier updates, (2) algorithmic changes to the estimator/objective/baseline, and (3) system-side changes that reduce sampler–learner mismatch.
@@ 2.4 Asynchronous distributed RL
3 VCPO: Robust, Scalable Asynchronous Post training
3.1 Why does Asynchronous Off-Policy RL collapse?
@@
Baseline choice for diagnostics. To isolate the effect of policy lag, we adopt a strong masking/clipping baseline for these didactic collapse experiments. Based on the masking and clipping sweeps in Appendix E.2, we find that sequence-level truncated importance sampling (TIS) with a higher threshold ($c=8.0$) delays collapse the longest among masking/clipping methods. We therefore use this setting as our default off-policy correction when diagnosing collapse.
Optimizer and KL regularization. Unless otherwise noted, we use AdamW with weight decay (see Appendix F.1 for details). Following prior asynchronous Fu et al. (2025); Wu et al. (2025); Espeholt et al. (2018) and synchronous Yu and others (2025) RL works, we do not incorporate explicit KL regularization loss as it constrains learning ability. We corroborate this observation with our own experiments (Appendix E.3).
@@ 3.3 Minimum Variance Baselines for Off-Policy RL
Minimum Variance Off-policy Baseline. As we derive in Appendix B), minimizing $\mathrm{Var}(\hat{G}(b))$ over scalar $b$ yields a closed-form solution for the off-policy optimal baseline (OPOB)
@@ 3.4 Efficient Baseline-Aware Gradient Computation
$\hat{G}(b)\;=\;\underbrace{\frac{1}{B}\sum_{i=1}^{B}w_{i}R_{i}g_{i}}_{\text{reward-weighted term}}\;-\;b\cdot\underbrace{\frac{1}{B}\sum_{i=1}^{B}w_{i}g_{i}}_{\text{score term}}.$ (8)
Thus, for any chosen $b$, the final gradient can be formed by combining two aggregated quantities:
$S_{R}\;=\;\sum_{i=1}^{B}w_{i}R_{i}g_{i},\quad S_{1}\;=\;\sum_{i=1}^{B}w_{i}g_{i}$
$\hat{G}(b)=\frac{1}{B}(S_{R}-bS_{S}).$
Score-gradient reweighting into two buffers. We take advantage of this linearity by computing per-sample score gradients $g_{i}$ once and accumulate them into two gradient buffers: (i) $G_{R}$ a buffer that accumulates the reward/advantage-weighted score gradients $R_{i}w_{i}g_{i}$, and (ii) $G_{S}$ a score buffer that accumulates score gradients $w_{i}g_{i}$. Afterwards, we form the final update by a simple linear combination of these buffers:
$G_{R}=\sum_{i}w_{i}R_{i}g_{i},\quad G_{S}=\sum_{i}w_{i}g_{i}.$
$\Rightarrow\hat{G}(b)=\frac{1}{B}\left(G_{R}-b\,G_{S}\right),\quad$
This avoids any additional backward pass to incorporate the baseline term.
@@ 3.5 VCPO: Variance Controlled Policy Optimization
As noted in Appendix E.2, we find that sequence-level truncated importance sampling (TIS) with a higher threshold ($c=8.0$) delays collapse the longest among masking/clipping methods. We thus combine sequence-level TIS with our effective-sample-size step-size scaling and our off-policy optimal baseline. This leads to the following surrogate loss, whose gradient equals our baseline-aware, truncated IS policy-gradient estimator
@@ 4 Empirical Evaluation
Figure 5: Qwen2.5-7B under PipelineRL-10 (10 Steps Off-Policy) on Countdown and MATH-500. Across both tasks, sequence level truncated importance sampling (TIS) suffers ESS-ratio collapse followed by KL/gradient instability and degraded accuracy, whereas VCPO maintains healthy ESS and stable updates, reaching synchronous performance. Training details and hyperparameters are provided in Appendix F.3.
@@ 4.1 Tasks
We evaluate VCPO across three representative RL post-training pipelines for LLM reasoning and agentic tool use. These settings require long rollouts (reasoning and multi-turn interactions with tool calls), so training is often dominated by rollout generation. As a result, strictly on-policy training underutilizes compute: the gradient updates must wait for sampling to complete, and this bottleneck worsens at scale and in sparse-reward domains that require more exploration.
We test VCPO in highly asynchronous regimes. Across all three pipelines, VCPO remains robust to substantial learner–sampler mismatch, improving throughput while preserving strong learning performance
@@
    Tool-integrated multi-turn reasoning. We evaluate multi-turn tool use in the SimpleTIR setting Xue et al. (2025), where the model must interleave reasoning with external tool calls. We train using the DAPO dataset Yu and others (2025) and evaluate on a held-out exam-style benchmark (AIME2025).
@@ 4.2 Metrics and Baselines
Other Baselines. We compare against representative prior approaches spanning (i) masking/clipping-based stabilization (e.g., M2PO, geometric sequence masking, TIS), (ii) objective-level approaches such as GSPO and OTB, and (iii) numerical/systems mitigations (e.g. using fp16 precision). In this paper, we train dense models, and so MoE specific methods such as Rollout Routing Replay are not considered.
4.3 VCPO Matches Synchronous Performance with Better Throughput
Across mathematical reasoning tasks, VCPO matches the final performance of fully synchronous RL even under high asynchrony. On GSM8K and MATH-500, Figures 4 and 5 show that VCPO remains stable throughout training and reaches accuracy comparable to the synchronous reference, whereas several baselines collapse or crash in this regime. In a general reasoning task such as Countdown, Figure 5 shows that VCPO attains the same peak performance as synchronous RL with higher end-to-end throughput enabled by asynchronous training. Furthermore, component ablations show both ESS based learning rate scaling and the off-policy optimal baseline improve robustness, and combining them yields the best stability and accuracy (Appendix D).
@@ 4.4 VCPO Provides State-of-the-Art Stability Under High Asynchrony
In contrast, VCPO consistently avoids collapse. Across baselines: (i) ratio-based clipping (TIS) and masking (MIS) eventually collapse; (ii) algorithmic variants such as GSPO and OTB do not reliably prevent collapse; and (iii) systems-side mitigations (e.g., FP16) delay collapse but do not address the underlying instability mechanism. Although (ii) and (iii) were not designed specifically for highly asynchronous RL, we include them for completeness because they are commonly used to stabilize on-policy LLM-RL training.
@@
4.5 VCPO Enables Fast and Stable Long-Context Multi-Turn RL
@@
Due to longer trajectories and prevalence of out-of-distribution tool calls, we use a lower level of asynchrony (2 steps off-policy). Figure 1 shows that unlike sequence-level TIS which collapses, VCPO improves steadily in both train reward and AIME2025 accuracy, reducing compute by $2.5\times$ compared the synchronous baseline. This demonstrates that our variance-control mechanisms transfer to long-horizon, multi-turn post-training.
@@ 5 Discussion
Impact Statement
More stable asynchronous RL can reduce wasted compute and enable more efficient post-training for reasoning and tool use. However, improving reliability may also increase agent capabilities and potential misuse; deployments should therefore follow domain-appropriate safety evaluations and monitoring.
@@ Appendix A Why does the Asynchronous RL collapse?
Appendix B Optimal Off-Policy Reward Baseline
@@ F.2 GSM8K
We use PipelineRL-$k$ asynchronous rollouts with maximum policy lag k=12k{=}12, and compute sequence-level importance weights between the sampler and learner policies. All methods (including our baselines) share the same sampling/training pipeline and differ only in the variance-control and/or masking rules applied to the policy-gradient update.
@@ F.3 Countdown and MATH
For both tasks, we run PipelineRL-$k$ with maximum policy lag k=10k{=}10 and use sequence-level importance sampling for off-policy correction. We use $4\timesH$100 GPUs for sampling and $4\timesH$100 GPUs for training
@@ F.4 Multi-Turn Tool Integrated Reasoning
We follow the SimpleTIR setting Xue et al. (2025) for multi-turn tool-integrated reasoning. We train on the DAPO-17K dataset Yu and others (2025) and evaluate on held-out exam benchmarks (AIME 2024 and AIME 2025), reporting exact-match final-answer accuracy. We use Qwen2.5-7B Base (rather than newer Qwen3 variants) to reduce the risk of evaluation-set contamination.

### v2

@@ 1 Introduction
Policy-gradient (PG) methods directly optimize a parameterized policy from sampled trajectories and have become the dominant approach for improving reasoning in large language models. Common objectives include Proximal Policy Optimization (PPO) Schulman et al. (2017) as well as REINFORCE Williams (1992) and its recent variants such as GRPO Shao et al. (2024).
These methods are typically implemented in an on-policy or near on-policy regime, where data generation and learning must occur sequentially after each other. This coupling becomes a major systems bottleneck for long-horizon tasks: the distribution of rollout lengths exhibit a long tail Hu et al. (2025), leading to pipeline bubbles and poor GPU utilization in synchronous RL training.
[…]
In contrast, we identify a deeper optimization failure mode: asynchronous RL training does not explicitly account for the variance of the policy-gradient estimator itself. Across math, general reasoning, and tool-use benchmarks, we show that collapse is reliably predicted by a collapsing effective sample size (ESS) Kish (1995) of the importance-weighted gradient estimate, suggesting that collapse is preceded by increasing gradient variance.
Motivated by this insight, we propose Variance Controlled Policy Optimization (VCPO): a method with explicit variance-targeted controls that stabilize policy-gradient objectives for asynchronous RL training. VCPO consists of two complementary components: (i) an effective sample size guided step scaling that down-weights unreliable updates and (ii) a closed-form minimum-variance off-policy reward baseline based on per-trajectory gradient norms, avoiding training an auxiliary critic or value model. To make (ii) practical at scale, we present the first scalable implementation for exact per-example gradient-norm computation with negligible runtime overhead, and leverage these statistics to build a low-overhead, gradient-aware baseline for variance reduction.
@@ 2 Background
2.1 Reinforcement Learning for LLMs
@@
2.2 Why Effective Sample Size?
@@
2.3 Prior RL stabilization methods
While prior methods have been proposed to stabilize LLM RL in on-policy or near-on-policy settings, we discuss these works as they closely related to asynchronous training’s highly off-policy issues. These approaches broadly fall into three categories: (1) masking/clipping to suppress outlier updates, (2) algorithmic changes to the estimator/objective, and (3) system-side changes that reduce sampler–learner mismatch.
@@ 2.4 Asynchronous RL
3 VCPO: Robust, Scalable Asynchronous RL
We develop VCPO, a variance-controlled optimization framework for asynchronous RL. We first diagnose collapse through effective sample size (ESS), showing that dominated, high-variance updates precede KL and reward collapse. We then derive a ESS-aware learning rate scaling rule and the variance-minimizing off-policy baseline (OPOB) for importance-weighted policy gradients. Finally, we give a single-backward implementation and combine these components into VCPO.
3.1 Why does Asynchronous RL collapse?
@@
Baseline choice for diagnostics. To isolate the effect of policy lag, we adopt a strong masking/clipping baseline for these didactic collapse experiments. Based on the masking and clipping sweeps in Appendix E.2, we find that sequence-level truncated importance sampling (TIS) with a high threshold ($c=8.0$) delays collapse the longest among masking/clipping methods. We therefore use this setting as our default off-policy correction when diagnosing collapse.
Optimizer and KL regularization. For all experiments we use AdamW with weight decay (see Appendix F.1 for details). Following prior asynchronous Fu et al. (2025); Wu et al. (2025) and synchronous Yu and others (2025); Shah et al. (2026) RL works, we do not incorporate explicit KL regularization loss as it constrains learning ability (corroborated by our own experiments in Appendix E.3).
@@ 3.3 Minimum Variance Baselines for Off-Policy RL
Optimal Off-policy Baseline. As we derive in Appendix B), minimizing $\mathrm{Var}(\hat{G}(b))$ over scalar $b$ yields a closed-form solution for the off-policy optimal baseline (OPOB)
@@ 3.4 Efficient Baseline-Aware Gradient Computation
$\hat{G}(b)$ $=\frac{1}{B}\underbrace{\sum_{i=1}^{B}w_{i}R_{i}g_{i}}_{\text{reward-weighted term}}-b\cdot\frac{1}{B}\underbrace{\sum_{i=1}^{B}w_{i}g_{i}}_{\text{score term}}$ (8)
$=\frac{1}{B}\bigl(G_{R}-b\,G_{S}\bigr).$
where $G_{R}\triangleq\sum_{i=1}^{B}w_{i}R_{i}g_{i}$ and $G_{S}\triangleq\sum_{i=1}^{B}w_{i}g_{i}$. Thus, for any chosen $b$, the final gradient can be formed by combining two aggregated quantities $G_{R}$ and $G_{S}$.
Score-gradient reweighting into two buffers. We take advantage of this linearity by computing per-sample score gradients $g_{i}$ once and accumulate them into two gradient buffers: (i) $G_{R}$ a buffer that accumulates the reward/advantage-weighted score gradients $R_{i}w_{i}g_{i}$, and (ii) $G_{S}$ a score buffer that accumulates score gradients $w_{i}g_{i}$. Afterwards, we form the final update by a simple linear combination of these buffers following Eq. 8. This avoids any additional backward pass to incorporate the baseline term.
@@ 3.5 VCPO: Variance Controlled Policy Optimization
As noted in Appendix E.2, we find that sequence-level truncated importance sampling (TIS) with a higher threshold ($c=8.0$) delays collapse the longest among all current stabilization methods. We thus combine sequence-level TIS with our effective-sample-size step-size scaling and our off-policy optimal baseline. This leads to the following surrogate loss, whose gradient equals our baseline-aware, truncated IS policy-gradient estimator
@@ 4 Empirical Evaluation
Figure 5: Qwen2.5-7B under PipelineRL-10 (10 Steps Off-Policy) on Countdown and MATH-500. Across both tasks, sequence level truncated importance sampling (TIS) suffers ESS collapse followed by KL/gradient instability and degraded accuracy, whereas VCPO maintains healthy ESS and stable updates, reaching synchronous performance. Training details and hyperparameters are provided in Appendix F.3.
@@ 4.1 Tasks
We evaluate VCPO across three representative RL post-training pipelines for LLM reasoning and agentic tool use in highly asynchronous regimes. Across all three domains, VCPO remains robust to substantial learner–sampler mismatch, improving throughput while preserving strong learning performance
@@
    Long horizon tool-integrated reasoning. We evaluate multi-turn tool use in the SimpleTIR setting Xue et al. (2025), where the model must interleave reasoning with external tool calls. We train using the DAPO dataset Yu and others (2025) and evaluate on a held-out exam-style benchmark (AIME2025). This setting requires long rollouts (reasoning and multi-turn interactions with tool calls), so training is often dominated by rollout generation. As a result, strictly on-policy training underutilizes compute: the gradient updates must wait for sampling to complete, and this bottleneck worsens at scale and in sparse-reward domains that require more exploration.
@@ 4.2 Metrics and Baselines
Other Baselines. We compare against representative prior approaches spanning (i) masking/clipping-based stabilization (e.g., M2PO, MIS, TIS), (ii) objective-level modifications such as GSPO and OTB, and (iii) numerical/systems mitigations (e.g. using fp16 precision). Although (ii) and (iii) were not designed specifically for highly asynchronous RL, we include them for completeness because they are commonly used to stabilize on-policy LLM-RL training. In this paper, we train dense models, and so MoE specific methods such as Rollout Routing Replay are not considered.
4.3 VCPO Delivers Synchronous Performance with Higher Throughput
Across mathematical reasoning, general reasoning, and tool-use tasks, VCPO matches the final performance of fully synchronous RL even under high asynchrony. On GSM8K, Figures 4 show that VCPO enables stable training while several baselines collapse or crash in this regime. For the Countdown and MATH-500 tasks, Figure 5 shows that VCPO attains the same peak performance as synchronous RL with higher end-to-end throughput enabled by asynchronous training. Table 1 summarizes the end-to-end wall-clock gains of up to $2.5\times$. Furthermore, component ablations show both ESS based learning rate scaling and the off-policy optimal baseline improve robustness, and combining them yields the best stability and accuracy (Appendix D).
@@ 4.4 VCPO Provides State-of-the-Art Stability Under High Asynchrony
In contrast, VCPO consistently avoids collapse. Across baselines: (i) ratio-based clipping (TIS) and masking (MIS) eventually collapse; (ii) algorithmic variants such as GSPO and OTB do not reliably prevent collapse; and (iii) systems-side mitigations (e.g., FP16) delay collapse but do not address the underlying instability mechanism.
@@
Scaling to extreme staleness. While the end-to-end training speedups from asynchronous are fully saturated for $k=10$ steps off-policy, for completeness we stress test stability as policy lag increases. In Figure 6, we see VCPO training remains stable up to at least k=128k{=}128 steps off-policy while matching synchronous training performance (k=0k{=}0).
4.5 VCPO Enables Fast and Stable Long-Horizon RL
@@
Due to longer trajectories and prevalence of out-of-distribution tool calls, we use a lower level of asynchrony (2 steps off-policy). Figure 1 shows that unlike sequence-level TIS which collapses, VCPO improves steadily in both train reward and AIME2025 accuracy, reducing end-to-end training time–a by $2.5\times$ compared the synchronous baseline. This demonstrates that our variance-control mechanisms transfer to long-horizon, multi-turn post-training.
Figure 6: Stable training under extreme policy lag: Qwen2.5-7B trained with PipelineRL at policy lag $k$ (steps off-policy). VCPO is stable up to at least k=128k{=}128 steps off-policy and matches the synchronous baseline (k=0k{=}0).
@@ 5 Discussion
Acknowledgements
We thank MIT-IBM Watson AI Lab, Amazon, Dell, and National Science Foundation for supporting this research. We thank NVIDIA for donating the DGX server.
@@ Appendix A Why does the Asynchronous RL collapse?
Appendix B Off-Policy Optimal Baseline
@@ F.2 GSM8K
[…]
Note on Hardware.
Prior work Liu et al. (2025a) reports that Ampere-generation GPUs can exhibit larger sampler–learner numerical discrepancies (e.g., in bf16 log-prob computations) which are not observed on Hopper-generation hardware. This leads to training collapse even for fully on-policy RL without policy-lag. We intentionally run GSM8K on Ampere (A6000) to stress-test robustness under amplified mismatch in addition to policy lag. Accordingly, this serves as a worst-case stress test: absolute metrics may change on other hardware, but the relative stability gains of our method over baselines are consistent in our checks.
@@ F.3 Countdown and MATH
For both tasks, we run PipelineRL-$k$ with maximum policy lag k=10k{=}10. We use $4\timesH$100 GPUs for sampling and $4\timesH$100 GPUs for training.
@@ F.4 Multi-Turn Tool Integrated Reasoning
We follow the SimpleTIR setting Xue et al. (2025) for multi-turn tool-integrated reasoning. We train on the DAPO-17K dataset Yu and others (2025) and evaluate on held-out exam benchmarks (AIME 2025), reporting exact-match final-answer accuracy. We use Qwen2.5-7B Base (rather than newer Qwen3 variants) to reduce the risk of evaluation-set contamination.